# hard problem of consciousness

### v2

@@ 4.3 Preconditions
There are preconditions for the existence of a causal identity. First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions, meaning we have a vocabulary of sufficient scale.
Second, there must be an incentive to construct it. Inference is only possible if some states are preferable to others. It is a value judgement. As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
@@ 4.4 Realising Lower Order States And Higher Order Meta Representations
The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions. Learning is not just about constructing policies but joining them together into weaker, more abstract policies. By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
This is formalised in definition 6. An organism $\mathfrak{o}$ has a set of policies it has learned. Each and every policy implies a $\mathfrak{v}$-task. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant. Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$. Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm). In particular, the organism must always act according to an interpretation, and some interpretations imply others. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
Figure 2: Illustration of how the generational hierarchy of tasks is a lattice of protosymbols.
The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
@@ 5.1 The First Order Self
A first order self (1ST henceforth, see figure 3) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [merker2007, barron2016]. Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [pearl2018].
A 1ST order self formally defined in definition 13. It serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. But also, to anticipate future planned interactions (for example, an insect navigating its environment [barron2016]), and recollect past interactions (it is a subset of all those relevant “statements” in the formal language).
@@
Natural selection prefers efficiency. The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects.. Decentralised or asynchronous control might be advantageous in some circumstances (e.g. an ant colony), but not others (e.g. in an individual human body). An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”. Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [bennett2023b].
If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must learn one. This might be partly hard-coded, but our focus is on inference by the organism itself.
@@ 5.2 The Second Order Selves
More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^17] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@ 5.3 The Third Order Selves
We can continue scaling the vocabulary and capacity for WPO indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that self-organizing systems such as a human body processes information to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
@@
Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep). To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena. All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
@@
There would, however, be a very clear delineation between conscious and not. The absence of a 1ST order self would mean there would be no one policy we might point at to say “this is what it is like” to be an organism. No policy linking all interventions together, and so no “self” that experiences making them. Hence a 1ST order self must precede all others.
@@ 7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from contentless physical states, up to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self, as does human behavior [^18] We additionally speculate that different orders of learned self may interact through hierarchical planning in a distributed system, perhaps something like a connective core [shanahan2012]. Arguably [boltuc2012], the full richness of human consciousness may be distinguished from phenomenal and access consciousness alone..
Conscious “access” as a human has it involves meaningful report in the sense described by Grice, which as we’ve established requires 2ND order selves. This is very different from the mere “access” to information. Hence, we hold that contents of which one is “access conscious”, in the sense that they are available for reasoning and report, can only be the contents of 2ND order selves. However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning. For that, one requires a 1ST order self. Hence access consciousness requires both 1ST and 2ND order selves.
Hence our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is impossible. For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational. In pancomputational enactivism there are only contentless states and the programs they imply. Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence. We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
@@
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [ball2023]. A human is a “multiscale competency architecture” [fields2020, levin2021, levin2024]. Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
@@ 8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 6). For each stage we point out animals which are likely to be at least so conscious:
$0:$ Unconscious (e.g. a rock)
@@
Stage 0: Unconsciousness
Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned.
  • Example: A rock.
@@
Stage one refers to adaptations hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [heylighen2023b] in a reasonably consistent environment.
@@
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by the nematode C. elegans [rankin2022, willett2018], which has a centralised nervous system and exhibits some ability to adapt with experience. However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978]. A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
@@
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [merker2005, merker2007, barron2016]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^19] In the same way declarative and imperative programs are equivalent [howard1980].. The “weaker” a policy is, the more behaviours it motivates. Conversely, if one’s actions share anything in common it is the intent that motivated them. The more diverse the actions, the more “general” or “high level” the intent they share. For example, the action of eating tends to involve the intent of satisfying hunger. If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour. While this is an unusual form of words, we contend that an organism is such a policy.
A stage three there is a self to feel simple hunger and pain. However, communication as described by Grice would be impossible [grice1969], because a stage three organism has no self-awareness, to represent another’s perception of their intent [bennett2023d]. Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
@@
Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. In other words, access consciousness follows phenomenal consciousness. The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
@@
  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. It allows one to anticipate how others will perceive one’s behaviour. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
@@
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. The seemingly altruistic behavior observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.

### v3

@@ 4.2 Preconditions
There are preconditions for the existence of a causal identity. First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be of sufficient scale to ensure that observations are distinguishable from interventions.
Second, there must be an incentive to construct it. Inference is only possible if some states are preferable to others. One cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
@@ 4.3 Realising Lower Order States And Higher Order Meta Representations
Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant. We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions. This serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [merker2007, barron2016], observed in mammals and insects.
@@
A 1ST order self may require centralisation and a “solid brain” with a persistent structure [sole2019] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
@@ 5.2 The Second Order Selves
More formally using the notation given in the appendix [bennett2024appendices] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@ 5.3 The Third Order Selves
We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that human body self organises to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
@@
Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep). To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena. All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
@@
There would, however, be a very clear delineation between conscious and not. The absence of a 1ST order self would mean there would be no policy linking all interventions together, and so no “self” that experiences making them. Hence a 1ST order self must precede higher orders.
@@ 7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [boltuc2012] depends upon the interaction of selves, perhaps in planning [shanahan2012]..
Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves. This is very different from the mere “access” to information. However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning. For that, one requires a 1ST order self. Hence access consciousness requires both 1ST and 2ND order selves.
Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational. In pancomputational enactivism there are only contentless states and the programs they imply. Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence. We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
@@
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [ball2023], and human intelligence is the high level goal directed behaviour of a swarm of cells [fields2020, levin2021, levin2024]. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
@@ 8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for WPO, through natural selection. For each stage we point out animals which are likely to be at least so conscious:
$0:$ Inert (e.g. a rock)
@@
Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [heylighen2023b] in a stable environment.
@@
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by nematodes [rankin2022, willett2018] with a centralised nervous system and some ability to adapt with experience. However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978].
@@
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [merker2005, merker2007, barron2016]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [howard1980].. For example, eating tends to involve the intent of satisfying hunger.
A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective. Subsequently it cannot communicate in the Gricean sense [grice1969, bennett2023d], or conceive of its own death, or experience shame.
@@
Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
@@
  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
@@
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. Altruistic behaviour observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.